Crew 222 Photos February 27th

Preview attachment 0520025_4251595_27Feb2020_Aventuriers.jpeg0?ui=2&ik=d401e10b9d&attid=0.1&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ-hK7_Dv_yReSOq4T9NQzZPAvkaeoqQAMSosGshb_QcWrmSh_uh9pBbQVbOrxNZ9Q5vaLjQBI7gU-8KV-KjBJrs1vNWJ6_35ugL7xoUuk4aZdjKVUpktYnHC2g&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
0520025_4251595_27Feb2020_Aventuriers.jpeg
288 KB

Preview attachment 0520025_4251595_27Feb2020_Canyon Walls.jpeg0?ui=2&ik=d401e10b9d&attid=0.2&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ9eKrecXYI1jtMulvZ5ro9GyLIkdDdgrcoyd1MXPp_OA9RVVygJbZWl-t4yheT6WaEnR7MhhH9Vd4XAMJjn16VE8LIMRTyEx_eIpM_Uv-itAZkLY4fxAkU-avo&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
0520025_4251595_27Feb2020_Canyon Walls.jpeg
368 KB

Preview attachment 0520283_4251199_27Feb2020_Hide and Seek.jpeg0?ui=2&ik=d401e10b9d&attid=0.6&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ8I7Je9zkiQa4Scb5B_JzDVHhB1jYxdG_nPx51JvD3f0ho3grpWojJdZCQ46u1oFroCiB2xsPNFJXwf1WOiGrLqqrBXgscoTdmlh01DsQXzic-08ENFOstpXng&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
0520283_4251199_27Feb2020_Hide and Seek.jpeg
389 KB


Preview attachment 0520283_4251199_27Feb2020_Face to face with time.jpeg0?ui=2&ik=d401e10b9d&attid=0.5&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ-yxJ9ROiTeFcQWJJWSSeUmyqUDfZdFTHRo-WWi2DHsTiuhtlckCKr-p76ONojdSIJ-qQrCuy1jfpnRnQI-khIWNosndZBezojZM9rDES3Ct3GusKmWWPjxYvY&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
0520283_4251199_27Feb2020_Face to face with time.jpeg
347 KB

Preview attachment 0520025_4251595_27Feb2020_Map Reading.jpeg0?ui=2&ik=d401e10b9d&attid=0.3&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ_VTMUlprRQ0WZeM3dIbymQZHthvHurYeWyFi-t8fjwCrWXYRG9AQsdKZw3YCcvA-RCpISxWydZ1b-HXyeUXJbYBDAGihsBj_zoKBonLl0MElvNYKGPavStIBs&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
0520025_4251595_27Feb2020_Map Reading.jpeg
259 KB

Preview attachment 0520283_4251199_27Feb2020_Canyon Walk.jpeg0?ui=2&ik=d401e10b9d&attid=0.4&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ_JKD2ISFqHS8CAFjySbgyGubyOVHRCj8CVWRJI8lHKafEUpknrJYDF62ceD7niMhthpi9zFfDq86leQTiqzPUV4_z_FRtevDZm4XjUzOyijpDls-9wo59DSgQ&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
0520283_4251199_27Feb2020_Canyon Walk.jpeg
335 KB

Preview attachment Science Dome_27Feb2020_AeroSpring.jpeg0?ui=2&ik=d401e10b9d&attid=0.7&permmsgid=msg-f%3A1659748368146513530&th=17089c58dbe83e7a&view=fimg&disp=thd&attbid=ANGjdJ8DJOm2OkIxVvKWmg6iu4wrwmX6pHsFZ99vWEygC6apV_CM2_1Pbfur2I7HndNjFRfvMOfep_0IGPa-RMcyLBylqNvh5o1LXkZuhlxIUqEmpclPQgJctlQZX6Y&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
Science Dome_27Feb2020_AeroSpring.jpeg
72 KB

Crew 222 Astronomy Report 27Feb2020

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/27/2020

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one) MDRS-WF

Objects to be Imaged this Evening: The pending observation of M43 should
be completed tonight.

Images submitted with this report: The final image of M51 was rendered
today and is attached to this report.

Problems Encountered: The red filter seems to be particularly affected
by the bad pixels issue. However, this has been mitigated so far by
using the Dust&scratches tool in Photoshop.

MUSK OBSERVATORY

Solar Features Observed: The observatory was used today for solar
imaging, but no interesting feature is to be reported.

Images submitted with this report: None

Problems Encountered: None

Preview attachment M51_200226.jpg0?ui=2&ik=d401e10b9d&attid=0.1&permmsgid=msg-f%3A1659747491354925791&th=17089b8cb70e5adf&view=fimg&disp=thd&attbid=ANGjdJ9AnYN0-COhbHfmgUAZRe6-8DG34fNd36bK45kyTugRa-Z26t9L-S8lNWe3Rl2M0DPmlpfWPVOBQSNeYM3P3zDnnJRT3fWnvZJMcMwjG5B7nIjnQakUn5uIPmg&ats=2524608000000&sz=w180-h120-df-p-nucleardot.gif
icon_1_image_x16.png
M51_200226.jpg
31 KB

Crew 222 EVA Report February 27th

Crew 222 EVA Report 27-02-2020
EVA #12
Author: Benjamin Auzou (CO)
Purpose of EVA: Replacement of the batteries of MegaAres and LOAC. Then
exploration from Galileo road to Candor Chasma.
Start time: 9:08am
End time: 12:30am

Narrative: We quickly replaced the batteries of MegaAres and LOAC. We
then headed towards Galileo road with 3 ATVs and Deimos. On the road, we
met a group of antelopes that crossed the road in front of us, we
managed to drive slowly and tried to disturb them as less as possible.
We stopped at the end of the road and then walked toward the canyon. We
walked along it and stopped at 11am to turn back. The road back to the
Hab was nominal. Near the Hab we took five minutes to take portraits for
the communication of our mission.

Destination: Experiments site and Candor Chasma
Coordinates (use UTM NAD27 CONUS): 12S 0518494-4250679 and 12S
0519900-4251500
Participants: Benjamin Auzou (CO, EVA leader), Bleuenn Roiron (GHO),
Marie Bochard (Journalist), Nicolas Bourdeaud (HSO)
Road(s) and routes per MDRS Map: Cow Dung Road, Hab Ridge Road
Mode of travel: Deimos and ATVs 1, 2 and Honda

Crew 222 GreenHab Report February 27th

Crew 222 GreenHab Report – 27/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Both heating and ambiant, door open during
the afternoon

Average temperatures: 20,5°C

Hours of supplemental light: 19h30 PM to 00h

Daily water usage for crops: 12 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 35,5 gallons

Time(s) of watering crops: 2

Changes to crops: We planted cilantro and some sun carrots.

Narrative:This morning, I was in EVA so Nina took care of the GreenHab
and watered plants.
During the afternoon, I went to the GreenHab with our crew Commander
Benjamin clean the Greenhab. We harvested the last sun carrots and
planted some new ones, and cilantro.
Tonight i watered again plants, I also harvested some Lettuce, cilantro
and basil for our tonight vegetarian burgers! I also harvested mint for
a tea.

Harvest:
Sun Carrots:50g
Lettuce Buttercrunch:50g
Basil:15g
Lemon balm:2g
Mint:2g
Pinaple mint:2g
Cilantro:5g

Support/supplies needed: As the amount of water in the Water Tank is
unprecise, i don’t know if i’ll be able to water plants tomorrow with
the water in it, am i authorized to use the tap relied to the static tank?

Crew 222 Journalist Report February 27th

Crew 222 Journalist Report 27Feb2020

Author: Marie Bochard, Crew Journalist

Sol 11

Title: Candor Chasma’s Charisma

For this morning’s EVA, we headed out towards Candor Chasma and the surrounding area by ATV. As scary as it is driving these beasts, the drive was fun and we all arrived safely at the end of the road. We even saw some antelopes that crossed the path in front of us! There were about 7 of them fleeing from the noisy quads. It’s strange seeing another form of life here, in this region considered to be so hostile to life! We walked down to the canyon, with the high snowy mountains in the background.

It was a beautiful walk through the canyon. We could see layers of history on either side of us, layers of colored stone forged by the ages. Red, white, green, all the colors were visible one after the other. We walked until the time forced us to stop, soaking in the beautiful canyon walls and the traces of antelopes on the path.

The walk back was rather exhausting as we had to climb back the way we came down, so we were glad to be back in the Hab. Riding the ATVs was definitely a fun part of the stay here.

This afternoon, Nina finished setting up the AeroSpring tower. The seeds planted last week in the moss grew enough to be transferred in the culture tower. What an exciting day. We filled the water tank below the tower with water, so that the pump can spray water inside the tower and water the plants. Lots of water involved. We turned the power on and hurray! It works. She will now come twice a day to turn on the pump that waters the plants, in hope that they will grow big enough to be harvested by the next crew!

Before dinner, we decided to play cards for our last evening together in the sim. We discovered new games and new rules, and had a lot of fun. Tonight, for dinner, the menu is red bean burger with chili and homemade bread! Spicy and delicious.

Crew 222 Operations Report-February 27th

Crew 222 Operations Report 02/27/2020

SOL: 11

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Tuesday generator:
Off at 6:00am
On at 9:00pm SOC was 85%

Wednesday generator:
Off at 5:30am
On at 9:00pm SOC was XX%

Diesel Reading – 72%

Station Propane Reading – 84% (delivered this morning)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 200 gallons

Water Meter: 149548.32 units

Water (static tank) – 218 gallons

Static to Loft Pump used – yes

Water in GreenHab – 35.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (also moved by Dr Rupert)
Hours: 146.7
Beginning charge: 100%
Ending charge: 89%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.8
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: no
Hours: 138.8
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes (Honda, 1 and 2)
Reason for use: Run during the engineer EVA for the battery and for the EVA.
Oil Added? no
ATV Fuel Used: 20% of the tank.
Hours the ATVs were Used today: 30 minutes
Notes on ATVs: All ok!

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.
The suit number 7 was working well at the beginning but two hours after
the beginning of the EVA the fan was stopping at random moments but was
restarting after. The fan never stopped totally but wasn’t working
optimally.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Sol Summary Report – February 27th

Crew 222 Sol Summary Report 27-02-2020
Sol: 11
Summary Title: A calm Hab
Author’s name: Benjamin Auzou
Mission Status: all things running, and crew transition ready

Sol Activity Summary: After the matinal workout session, four of us went
on EVA to explore Candor Chasma, that was the last EVA for three of us.
In the afternoon, we worked on different activities : installation of
seeds in the culture tower, GreenHab cleaning, human factors experiments !
At 4pm, our HSO Nicolas leaved the Hab to Grand Junction and will be
back tomorrow afternoon.

Look Ahead Plan: Tomorrow will be the last day of sim, we will prepare
the arrival of our friends of the crew 223. We will also take time to
clean the station. The last EVA team of our rotation will explore the
area around Copernicus Highway.

Anomalies in work: none
Weather: sunny
Crew Physical Status: all members safe and sound
EVA: Exploration of Candor Chasma
Reports to be filed: Journalist report, Operations report, GreenHab
Report, Astronomy Report, EVA Report, EVA Request, Photos Report, Photo
of the day
Support Requested: nothing to report

Sol Summary – February 26th

Crew 222 Sol Summary Report 26-02-2020
Sol: 10
Summary Title: An other martian day
Author’s name: Benjamin Auzou
Mission Status: all things runnning, preparing the end of the rotation

Sol Activity Summary: In the morning, the crew woke up with a matinal workout session. Then four of us went on EVA to explore Hab Ridge, this. That was a calm and sunny day on Mars. In the afternoon we performed the bomb defusal cooperation game for the third time. Before the end of the afternoon, the HSO and the engineer get outside to take care of one of the ATVs. For dinner, we will taste the wonderful cucumber of the greenhab we are talking about for some days now.

Look Ahead Plan: Tomorrow, we will have our last real martian day together, because one of our crewmembers will leave tomorrow evening to get back to Earth.
We will continue the experiments and prepare the rotation with crew 223.

Anomalies in work: none
Weather: sunny
Crew Physical Status: all members safe and sound
EVA: Exploration of Hab Ridge
Reports to be filed: Journalist report, Operations report, GreenHab Report, Astronomy Report, EVA Report, EVA Request, Photos Report, Photo of the day
Support Requested: nothing to report

Operations Report – February 26th

Crew 222 Operations Report 02/26/2020

SOL: 10

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Tuesday generator:
Off at 5:30am
On at 8:45pm SOC was 80%

Wednesday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 72%

Station Propane Reading – 84% (delivered this morning)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149505.59 units

Water (static tank) – 258 gallons

Static to Loft Pump used – yes

Water in GreenHab – 47.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (also moved by Dr Rupert)
Hours: 146.3
Beginning charge: 100
Ending charge: 97%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.8
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.8
Beginning charge: 100%
Ending charge: 94%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery. Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The red one was jumped bu Dr Rupert during the afternoon. One of the blue ones doesn’t start. We will try to charge it during the night thanks to the battery charger of the RAM.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA. The suit number 7 was working well at the beginning but one hour after the beginning of the EVA the fan was stopping at random moments but was restarting after. The fan never stopped totally but wasn’t working optimally.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A